Two equally charged small balls are 3cm apart I'm air and repel each other with a force of 40μN.Compute the charge of each Ball.


1
Expert's answer
2021-04-01T19:00:20-0400
"F=\dfrac{kq^2}{r^2},""q=\sqrt{\dfrac{Fr^2}{k}}=\sqrt{\dfrac{40\cdot10^{-6}\ N\cdot(0.03\ m)^2}{9\cdot10^9\ \dfrac{Nm^2}{C^2}}}=2\cdot10^{-9}\ C=2\ nC,""q_1=q_2=q=2\ nC."